# Exploring the Potential of QEEGNet for Cross-Task and Cross-Dataset Electroencephalography Encoding with Quantum Machine Learning

### v1

@@ 2 Methodology
Figure 2: Structure of the Variational Quantum Circuit (VQC). The input $\vec{x}$ represents EEG features encoded by a classical neural network before being processed by the quantum circuit. The circuit consists of an encoding layer $U(\vec{x})$, variational layers $V_{i}(\vec{\theta}_{i})$, and measurement operations to extract quantum outputs.
@@ 2.2 Quantum Encoding Layer
$f_{\text{HNN}}:\mathbb{R}^{d}\to\mathbb{R},\quad f_{\text{HNN}}=M\circ U\circ V\circ C$ (3)
where $C:\mathbb{R}^{d}\to\mathbb{R}^{d^{\prime}}$ is the classical neural network, mapping input data $x\in\mathbb{R}^{d}$ to a lower-dimensional embedding space. $V:\mathbb{R}^{d^{\prime}}\to\mathcal{U}(\mathcal{H})$ is the encoding function that transforms classical features into quantum states. $U:\mathcal{U}(\mathcal{H})\to\mathcal{U}(\mathcal{H})$ is the variational quantum circuit responsible for parameterized quantum transformations. $M:\mathcal{H}\to\mathbb{R}$ is the measurement function that collapses the quantum state into classical output.
The computation process follows:
$f_{\text{HNN}}(x)=\langle\psi_{0}|V^{\dagger}(C(x))U^{\dagger}(\theta)HU(\theta)V(C(x))|\psi_{0}\rangle$ (4)
where $V^{\dagger}$ and $U^{\dagger}$ denote the adjoint operators of the encoding and variational quantum circuit, respectively.
The classical neural network is composed of multiple layers:
@@
where $W_{i}\in\mathbb{R}^{d_{i}\times d_{i-1}}$ and $b_{i}\in\mathbb{R}^{d_{i}}$ are the weight matrices and biases of the neural network, and $\sigma$ is a non-linear activation function (e.g., ReLU or Sigmoid). The output $C(x)\in\mathbb{R}^{d^{\prime}}$ serves as input to the quantum circuit.
The variational quantum circuit consists of a sequence of parameterized unitary operations:
$U(\theta)=U_{L}(\theta_{L})\cdots U_{1}(\theta_{1}),\quad U_{j}(\theta_{j})=e^{-i\theta_{j}H_{j}}$ (6)
where $H_{j}$ is a Hermitian operator defining the quantum transformation.
Finally, measurement is performed using a Hermitian observable $H$:
$M(|\psi\rangle)=\langle\psi|H|\psi\rangle$ (7)
The overall optimization is performed by minimizing the loss function:
$\mathcal{L}(\theta)=\sum_{i=1}^{N}\left(f_{\text{HNN}}(x^{(i)})-y^{(i)}\right)^{2}$ (8)
@@
Once EEG features are extracted by the classical neural network, they are transformed into quantum representations through a series of quantum encoding operations. Each extracted feature vector $z\in\mathbb{R}^{d}$ is mapped onto a quantum state via a parameterized rotation operation like Figure 3.
@@
The encoding utilizes a rotation-$Y$ (RY) gate:
$R_{Y}(z_{i})=\exp\left(-iz_{i}\frac{\sigma_{y}}{2}\right),$ (9)
where $\sigma_{y}$ is the Pauli-$Y$ matrix:
$\sigma_{y}=\begin{pmatrix}0&-i\\ i&0\end{pmatrix}.$ (10)
This transformation embeds classical data into a quantum Hilbert space, enabling richer representations.
@@
${CNOT}=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\end{pmatrix}.$ (11)
@@
$CNOT(q_{i},q_{(i+1)\mod n_{\text{qubits}}})\quad\text{for}\quad i=0,1,\ldots,n_{\text{qubits}}-1.$ (12)
@@ 2.3 Measurement and Classification
After quantum transformations, qubit states are measured to extract classical information for downstream classification. The expectation value of the Pauli-$Z$ operator is computed for each qubit:
$\langle\sigma_{z}^{i}\rangle=\langle 0|U^{\dagger}\sigma_{z}^{i}U|0\rangle,$ (14)
where $U$ represents the full unitary evolution of the quantum circuit. The Pauli-$Z$ operator, also known as the Pauli-$Z$ matrix, is represented by the following 2x2 matrix:
$\sigma_{z}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}.$ (15)
The measurement process collapses quantum states into classical probabilities, forming a feature vector suitable for final classification.
These extracted classical features are subsequently passed into a fully connected neural network layer. The final classification output is computed using the softmax function:
$P(y=j|x)=\frac{\exp(a_{j})}{\sum_{k}\exp(a_{k})},$ (16)
where $a_{j}$ represents the logits for class $j$. This final step produces the probability distribution over output classes, enabling accurate EEG-based categorization.

### v2

@@ 2 Methodology
Figure 2: Structure of the Variational Quantum Circuit (VQC). The input $\vec{x}$ represents EEG features encoded by a classical neural network before being processed by the quantum circuit. The circuit consists of an encoding layer $U(C(\vec{x}))$, variational layers $V_{i}(\vec{\theta}_{i})$, and measurement operations to extract quantum outputs.
@@ 2.2 Quantum Encoding Layer
$f_{\text{HNN}}:\mathbb{R}^{d}\to\mathbb{R},\quad f_{\text{HNN}}=M\circ V\circ U\circ C$ (3)
where $C:\mathbb{R}^{d}\to\mathbb{R}^{d^{\prime}}$ is the classical neural network, mapping input data $x\in\mathbb{R}^{d}$ to a lower-dimensional embedding space with $d>d^{\prime}$. A quantum embedding unitary $U:\mathbb{R}^{d^{\prime}}\to\mathcal{U}(\mathcal{H})$ is defined to transform an embedded classical input $C(x)$ into a quantum state by |$\ket{\psi}\mapsto U(C(x))\ket{\psi}$. Subsequently, we define a variational quantum circuit $V:\mathcal{U}(\mathcal{H})\to\mathcal{U}(\mathcal{H})$ with tunable parameters $\theta$ to generate a corresponding quantum transformation $V(\theta)$. A final step of VQC is to perform a measurement on the resulting quantum state with a selected observable $M:\mathcal{H}\to\mathbb{R}$. The computation process above combined yields,
$f_{\text{HNN}}(x)=\langle\psi_{0}|U^{\dagger}(C(x))V^{\dagger}(\theta)HV(\theta)U(C(x))|\psi_{0}\rangle$ (4)
where $V^{\dagger}$ and $U^{\dagger}$ denote the Hermitian conjugates of the encoding and variational quantum circuit, respectively.
The classical embedding $C(x)$ is implemented by a classical neural network of multiple layers,
@@
where $W_{i}\in\mathbb{R}^{d_{i}\times d_{i-1}}$ and $b_{i}\in\mathbb{R}^{d_{i}}$ are the weight matrices and biases of the neural network, and $\sigma$ is a non-linear activation function (e.g., ReLU or Sigmoid). The output $C(x)\in\mathbb{R}^{d^{\prime}}$ serves as an input to the quantum circuit.
The variational quantum circuit consists of a sequence of parameterized unitary operations $V(\theta)=V_{L}(\theta_{L})\cdots V_{1}(\theta_{1})$ with $V_{\ell}(\theta_{\ell})=e^{-\frac{i}{2}\theta_{\ell}\sigma_{\ell}}$ for $\ell=1,\ldots,L$, and $\sigma_{\ell}$ is one of the Pauli matrices,
$\sigma_{x}=\begin{pmatrix}0&-1\\ 1&0\end{pmatrix},\quad\sigma_{y}=\begin{pmatrix}0&-i\\ i&0\end{pmatrix},\quad\sigma_{z}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}.$ (6)
depending on $\ell$. The measurement is performed by a selected Hermitian observable $H$ such that $M(|\psi\rangle)=\langle\psi|H|\psi\rangle\in\mathbb{R}$. A typical choice is the Pauli matrix in Eq. (6).
The optimization over tunable parameters $\theta$ is performed by minimizing the loss function:
$\mathcal{L}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\left(f_{\text{HNN}}(x^{(i)};\theta)-y^{(i)}\right)^{2}$ (7)
@@
Once EEG features are extracted by the classical neural network, they are transformed into quantum representations through a series of quantum encoding operations. Each extracted feature vector $z=(z_{1},\ldots,z_{d^{\prime}})\in\mathbb{R}^{d^{\prime}}$ is mapped onto a quantum state via a parameterized rotation operation like Figure 3.
@@
The encoding $U$ utilizes a rotation-$Y$ (RY) gate:
$R_{Y}(z)=\exp\left(-iz\frac{\sigma_{y}}{2}\right),$ (8)
where $\sigma_{y}$ is the Pauli-$Y$ matrix in Eq. (6). With this transformation, classical data is embedded into quantum states of a Hilbert space for richer representations.
@@
$\text{CNOT}=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\end{pmatrix}.$ (9)
@@
$\text{CNOT}(q_{i},q_{(i+1)\mod n_{\text{qubits}}})\quad\text{for}\quad i=0,1,\ldots,n_{\text{qubits}}-1.$ (10)
@@ 2.3 Measurement and Classification
After quantum transformations $U(x)$ and $V(\theta)$, qubit states are measured to extract classical information for downstream classification. Here, we choose a Pauli-Z matrix $\sigma_{z}$ in Eq. (6) so that the final VQC output in Eq. (4) yields $\braket{\sigma_{z}}=(y_{1},\ldots,y_{n})$.
The above measurement is then converted into classical probabilities via
$P(y=j|x)=\frac{\exp(y_{j})}{\sum_{k}\exp(y_{k})},\quad(j=1,\ldots,n)$ (12)
to form a vector for classification. The results are passed into a fully connected neural network layer, where the softmax function computes the final output, producing a probability distribution over output classes for accurate EEG-based categorization.